OpenStudy (anonymous):

you have 46 coins consisting of pennies, dimes and quarters. The number of pennies is equal to the number of dimes, and the total value of the coins is $5.26. How many of each denomination of coins do you have.

OpenStudy (radar):

Let x = then number of pennies and number of dimes (as they are equal) Then 46-2x would equal the number of quarters. Now to the value: x = the value of the pennies, 10x =value of dimes and 25(46-2x) = value of the quarters. This value sum amounts to $5.26 x +10x + 25(46-2x) = 526 11x + 1150 - 50x = 526 -39x=-624 x=16 16 pennies, 16 dimes 46-2x=46-32=14 14 quarters, number of coins =16+16+14 = 46 Good total value is 16 + 160 + 350 =526 cents or $5.26 That is good too.

OpenStudy (radar):

Did you follow with understanding?